How many atoms are in 11.5 moles of carbon?
Rounded to the appropriate number of Significant Figures. Give answer using scientific notation.
1
Expert's answer
2020-01-28T06:15:49-0500

The number of atoms in one mole of any substance is described by Avogadro's number (NA):

NA = 6.022 × 1023 atoms or particles in one mole

From here, the number of atoms in a substance equals:

N = n × NA

where N - the number of atoms, n - the number of moles, NA - Avogadro's number.

Therefore, as 1 mole of carbon contains 6.022 × 1023 atoms, 11.5 moles contain:

N = 11.5 mol × 6.022 × 1023 = 69.253 × 1023 = 6.93 × 1024 atoms of carbon


Answer: 6.93 × 1024